OpenStudy (anonymous):

Please Help! Solve the following system of equations and show all of your work. 4x + 3y = -1 3x + y = 3

OpenStudy (mathmate):

What methods have you worked on? For example, substitution, elimination, Cramer's Rule, comparison.

OpenStudy (anonymous):

elimination because its easier

OpenStudy (anonymous):

I got 4x + 3y = -1 3x + y = 3 x + 2y = 2 2y/2 = 2/2 y=1 but i dont know how to solve for x please help

OpenStudy (mathmate):

To solve by elimination, you need to multiply one or both equation by numbers so that the coefficient of one of the variables match. Add or subtract to eliminate the variable and subsequently solve for one of them. What you have done is to reduce the coefficient of x to 1, but not ELIMINATE. You are almost there. 4x + 3y = -1 ...(1) 3x + y = 3 ...(2) Multiply (2) by 3 and subtract (1), that gives you 3y-3y=0y so y is eliminated. The solve the remainder of the equation for x. To find y, use the value of x to substitute into equation (2) and solve for y.

OpenStudy (anonymous):

I kinda get it but can you show me the steps because i don't understand what you wrote?

OpenStudy (anonymous):

y = -4?

OpenStudy (mathmate):

4x + 3y = -1 ...(1) 3x + y = 3 ...(2) Multiply (2) by 3 and subtract (1) 9x + 3y = 9 ...(2a) 4x + 3y = -1 ...(1) subtract to get: 5x + 0y = 10 (3) Solve for x to get x=10/5=2 (4) Substitute (4) into (2) 3(5) + y = 3 ....(2) Can you solve for y now?

OpenStudy (anonymous):

4x + 3y = -1 ...(1) <------why do you put (2) ? 3x + y = 3 ...(2) <----(1)?

OpenStudy (mathmate):

The original equations are defined as: 4x + 3y = -1 ...(1) 3x + y = 3 ...(2) When we found the value of x in equation (4): x=10/5=2 (4) we have a choice of substituting x into equation (1) or (2).

OpenStudy (mathmate):

I have chosen to substitute (4) in (2) because the coefficient of y in (2) is 1, so there is less work: 3(2)+y=3 ....(2b) [ note, I have made a correction, since x=2 (not 5) in my previous solution] which makes 6+y=3 that solves to y=3-6=-3

OpenStudy (mathmate):

Now the only thing left to do is to substitute x=2 and y=-3 into the original equations (1) and (2) to verify that the solution is indeed correct.

OpenStudy (anonymous):

ohh ok but i did this 3 (4x + 3y = -1 ) -4 (3x + y = 3) 12x + 9y = -3 -12x - 4y = -12 5y = -15 y = -3

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

i get it now you have been a really great help(: thank you soo much!(:

OpenStudy (mathmate):

Yes, that's correct! You could eliminate either one. Eliminating x first gives you good practice!